OpenStudy (anonymous):

i read 4 questions per minute,i use 2 papers per question, how many papers are used in 1 hour.

OpenStudy (chihiroasleaf):

4 question per minute, and 2 papers per question, so how many paper do you use in one minute (for 4 questions) ?

OpenStudy (anonymous):

\[\frac{ 4 ? }{ \min }\frac{ 2p }{ ? }=\frac{ 8p }{ \min }\]

OpenStudy (anonymous):

4 per 1 2*4per 1min 8*60 per 60min

OpenStudy (anonymous):

4 questions/min * total number of min will give you total number od questions you can solve in that time(in this case it is 60min) . multiply this number 2 (number of papers you use per question) will finally give you number of papers you use in that time.

OpenStudy (anonymous):

i think 480 papers in 1 hour

OpenStudy (anonymous):

is the function of the papers P,and questions q, \[P(q)=q^2-\frac{ 2 }{ q }\] sorry this is correct

OpenStudy (anonymous):

and i read 4 questions per minute

OpenStudy (anonymous):

Find the rate at wich paper is used

OpenStudy (anonymous):

previous ? settled

OpenStudy (anonymous):

does that make sense

OpenStudy (anonymous):

\[\frac{ dP }{ dq }\frac{ dq }{ t }=\frac{ dP }{ dt }\]

OpenStudy (anonymous):

\[(2p+2q^{-2})(\frac{ dq }{ dt })=4\]

OpenStudy (anonymous):

\[\frac{ dq }{ dt }=\frac{ q^2 }{ 2q^3+2 }\] sorry only q

OpenStudy (anonymous):

yes now its a new question with a function

OpenStudy (anonymous):

so @theloser if it was cont the problem will be solved

OpenStudy (anonymous):

\[p'(t)=8q+8q^{-2}\]

OpenStudy (anonymous):

thanks guys i just made up the question s its not standard

OpenStudy (anonymous):

hey ok.. you can solve it now, dp/dt = (dp/dq)*(dq/dt)